OpenStudy (anonymous):

i really dont get how to do multi-step eqations... does anybody know and easy way to do this???

OpenStudy (anonymous):

It depends what equation your doing. What type of equation are you trying to figure out?

OpenStudy (anonymous):

something like -29=5(2a-1)+10

OpenStudy (goformit100):

Yes he/she is correct

zepdrix (zepdrix):

It's like a puzzle fenix, do you like puzzles? :D Err err it's more like.. those old waldo books, a = waldo. You gotta try to find out where he's hiding. You gotta move stuff around, check every corner! :O

OpenStudy (anonymous):

hmmmm.. yes i do like puzzles.... but how do you solve it?? my teacher explains in different ways... i get totally lost..

OpenStudy (anonymous):

sorry, its -29=5(2a-1-2a

zepdrix (zepdrix):

Your goal is to get "a" alone. Try to isolate it. So you'll notice there are a bunch of pieces around it, performing certain operations. > the 10 is being added to the side with a. Well one way to get closer to finding a, is to get the 10 on the side that DOESN'T have a. We're getting all the garbage off of a. So you want to perform the opposite operation to get the 10 over to the other side. Subtract 10 from both sides. It's not always that straight forward, but it's a start at least. <:o

zepdrix (zepdrix):

oh :D

OpenStudy (anonymous):

but ill use it... i got 15=5(2a-1)

zepdrix (zepdrix):